Three times the square of the number when added to seven times the number result in 26.What is the number ?
Answer by nerdybill(7384)

Three times the square of the number when added to seven times the number result in 26.What is the number ?
.
Let x = the number
.
3x^2 + 7x = 26
3x^2 + 7x - 26 = 0
(3x+13)(x-2) = 0
.
Setting each term to zero:
(3x+13) = 0
3x = -13
x = -13/3
.
(x-2) = 0
x = 2
.
Solution set:
x = {2, -13/3}
.
Tossing out the negative fraction leaves us with:
x = 2